Time for a lump sum to double

If you deposit money today in an account that pays 14.5% annual interest, how long will it take to double your money? Round your answer to two decimal places.
Answer #1

Let PV =100
FV =200
Rate =14.5%
FV=PV*(1+r)^n
200=100*(1+14.5%)^n
applying log on both sides we get =log(2)/log(1+14.5%) =5.12 years